# Maximum Length RLL Sequences in de Bruijn Graph

### v1

@@ Abstract
A timing and synchronization system based on a de Bruijn sequence has been proposed and studied recently for a channel associated with quantum communication that requires reliable synchronization. To avoid a long period of no-pulse in such a system on-off pulses are used to simulate a zero and on-on pulses are used to simulate a one. But, these sequences have high redundancy. To reduce the redundancy, run-length limited sequences in the de Bruijn graph are proposed for the same purpose. The maximum length of such sequences in the de Bruijn graph is studied and an efficient algorithm to construct a large set of these sequences is presented. A maximum length sequence for which the position of each window can be computed efficiently is constructed. Finally, an enumeration of the number of such sequences is given and some generalizations are discussed.
Index Terms: de Bruijn graph, necklaces, RLL sequences.
@@ 1 Introduction
In a free-space quantum key distribution, one of the goals is to synchronize the transmitted and received signals accurately. There are some efforts in designing efficient and reliable timing and synchronization systems [6, 18]. In [26] a de Bruijn sequence-based timing and synchronization system is introduced using a beacon with an on-off model. In this model, a sequence of beacon pulses is used to represent a binary de Bruijn sequence. Hence, once a sub-string of beacon pulses is received, its position is also determined uniquely. Furthermore, to consider the timing jitter performance, a long period of no-pulses is forbidden. Assume one pulse slot is used to represent a binary bit. If on-pulse is one and off-pulse is zero, then a long run of zeros in the sequence, which is a long period of no-pulses, would impact the timing jitter. In [26], two pulse slots are suggested to represent a single bit, on-on is a one and on-off is a zero so that two consecutive no-pulses are avoided. However, this scheme requires $2N$ pulse slots to represent a sequence of length $N$ in the order $n$ de Bruijn graph and it is required to receive a sub-sequence of $2n$ pulse slots to locate its position, i.e., if the sequence is long of about $2N=2^{n+1}$ bits (representing a sequence of length $N$ in the de Bruijn graph), about $2\log N$ pulse slots are required to locate its position instead of $\log N$ pulse slots if any sequence of $n$-pulses was permitted (all logarithms in this paper are in base 2). A second disadvantage in the proposed scheme is that the sequences of consecutive zeros were constrained to length one, while in reality, a few consecutive no-pulses are permitted subject to a constraint that their length will not be larger than a certain threshold $s$. Therefore, there is a target to use less redundant pulse slots to achieve both goals, to synchronize accurately, and to avoid long periods of no-pulses. For this purpose run-length limited (RLL) de Bruijn sequences are proposed [3, 4], where one pulse is represented by one binary bit. An on-pulse is represented by a one and an off-pulse is represented by a zero. This scheme is more general and more efficient with lower redundancy and higher rate than the ones in the previous work. The scheme combines two concepts, RLL sequences and sequences in the de Bruijn graph which will be defined now.
@@
RLL sequences are binary sequences in which there is an upper bound on the number of consecutive zeros in a sequence. An $(n,s)$-word is a binary word of length $n$ in which the longest run of consecutive zeros is of length at most $s$. An $(n,s)$-sequence is a sequence whose windows of length $n$ are distinct $(n,s)$-words. The family of sequences in which there are no runs of more than $s$ consecutive zeros was extensively studied due to many applications that require such sequences [16].
@@
The de Bruijn graph of order $n$, $G_{n}$, was defined first by Nicolaas Govert de Bruijn [1] and in parallel by Good [14]. The graph is a directed graph with $2^{n}$ vertices which are represented by the set of all binary words of length $n$. The edges of $G_{n}$ are represented by the $2^{n+1}$ binary words of length $n+1$. There is a directed edge $(x_{0},x_{1},\ldots,x_{n-1},x_{n})$ from the vertex $(x_{0},x_{1},\ldots,x_{n-1})$ to the vertex $(x_{1},\ldots,x_{n-1},x_{n})$. A span $n$ de Bruijn sequence is a cyclic binary sequence in which each binary $n$-tuple is contained exactly once in a window of length $n$. A walk (tour) in the graph is a sequence of directed edges such that the end-vertex of one edge is the start-vertex of the next edge. In a tour, the first vertex is also the last one. Any binary sequence can be represented by a walk in $G_{n}$, where any $n$ consecutive symbols represent a vertex and any $n+1$ consecutive symbols represent an edge. A cyclic sequence can be represented by a tour. The consecutive $n$ symbols and $n+1$ symbols represent a walk with its consecutive vertices and consecutive edges, respectively. A path (cycle) in the graph is a walk (tour) with no repeated vertices. In a cycle, each vertex can be considered the first vertex of the cycle. A trail (circuit) in the graph is a walk (tour) with no repeated edges. A span $n$ de Bruijn sequence can be represented by an Eulerian circuit in $G_{n-1}$, i.e., a circuit which traverses each edge exactly once. In a circuit, the first vertex is also the last one. It can be also represented by a Hamiltonian cycle in $G_{n}$, i.e., a cycle that visit each vertex of $G_{n}$ exactly once. Similarly, any path in $G_{n}$ can be represented by a sequence with no repeated $n$-tuples. This sequence can be represented also by a trail in $G_{n-1}$.
One structure that will be used in our exposition and is extensively studied in the literature associated with the de Bruijn graph is a necklace. A necklace of order $n$ is a cycle in $G_{n}$ whose length is a divisor of $n$. If the cycle is of length $n$, then the necklace is of full-order. If the necklace is of length which is a divisor of $n$ that is smaller than $n$, then the necklace is degenerated. Each vertex in $G_{n}$ is represented by a word of length $n$ and in the necklace, we have all cyclic shifts of such a word. The necklaces of order $n$ are the equivalence classes of the relation defined on words of length $n$, where two words are related if one is a cyclic shift of the other. If the necklace has length $d$ which is a divisor of $n$, then it contains $d$ words of length $n$, but the necklace can be represented by a word of length $d$. In $G_{n-1}$ such a necklace of order $n$ is also a cycle of length which is a divisor of $n$ with no repeated edges. The necklaces in $G_{n-1}$ is represented by the edges whose representation are words of length $n$. The runs of zeros in a necklace are considered to be a cyclic run since the necklace is a cycle. Let $(n,s)$-necklace be a necklace of order $n$ that does not have a (cyclic) run of more than $s$ consecutive zeros. Such a necklace contains only $(n,s)$-words.
@@
Assume $n=6$ and consider the $(6,2)$-necklace with the vertex $(001011)$. This necklace is of full-order and it contains the six words of length 6 in the cycle
@@
The degenerated necklace $[011]$ of order 6 contains the three vertices in $G_{6}$ which form the cycle of length 3 whose vertices and edge are given by
@@
In this work, we will consider a combination of sequences in the de Bruijn graph that have a run-length constraint on the number of consecutive zeros. We will be interested in the maximum length cyclic and acyclic $(n,s)$-sequences. By their definition, if $s$ is small, these sequences can be used in free-space quantum key distribution with flexibility in the parameters.
Let $G_{n}(s)$ be the subgraph of $G_{n}$ induced by all the vertices of $G_{n}$ whose representations are $(n,s)$-words. Our goal is to find the maximum length path and maximum length cycle in $G_{n}(s)$. The definition of $G_{n}(s)$ implies that the edges of this graph also do not have a run of zeros whose length is greater than $s$. This property will be important in $G_{n-1}(s)$. An upper bound on the maximum length of such a sequence will use a trail in $G_{n-1}(s)$ (which is a circuit if the sequence is cyclic) and a lower bound on the maximum length will use a path in $G_{n}(s)$ (which is a cycle if the sequence is cyclic). This is unique to our paper as usually it is not required to use different orders of the graph for similar tasks. Note, that a cyclic sequence of length $N$ with no repeated $n$-tuples has $N$ binary digits, $N$ vertices, and $N$ edges in $G_{n}$ and in $G_{n}(s)$. An acyclic sequence with no repeated $n$-tuples has $N+n-1$ binary digits, $N$ vertices, and $N-1$ edges in $G_{n}$ and $G_{n}(s)$. Henceforth, unless stated otherwise, all the $(n,s)$-sequences that will be considered will be cyclic.
@@
The following cycle in $G_{6}(2)$
$\rightarrow(100100)\rightarrow(001001)\rightarrow(010011)\rightarrow(100111)\rightarrow(001110)\rightarrow(011101)\rightarrow(111010)$
$\rightarrow(110100)\rightarrow(101001)\rightarrow(010010)\rightarrow(100101)\rightarrow(001011)\rightarrow(010111)\rightarrow(101110)$
$\rightarrow(011100)\rightarrow(111001)\rightarrow(110010)$
form the $(6,2)$-sequence
$10010011101001011$
of length 17. The sequence contains two words $(001110)$ and $(011100)$ which are $(6,2)$-words not in a $(6,2)$-necklace since their necklace contains the words $(000111)$, $(100011)$, $(110001)$, and $(111000)$, that are not $(6,2)$-words.
@@
Example 2 indicates that there could be long $(n,s)$-sequences which contain $(n,s)$-words that are not contained in $(n,s)$-necklaces. This implies that it is not straightforward to know the maximum length of an $(n,s)$-sequence.
@@
The rest of the paper is organized as follows. In Section 2 we consider an upper bound on the maximum length of $(n,s)$-sequences which are cycles in $G_{n}$. We prove that this bound is equal to the number of words in the $(n,s)$-necklaces. The trail with the maximum length can have another $s$ edges (and $s$ vertices) from $G_{n-1}(s)$. The proof will be based on the necklaces obtained by the edges of $G_{n-1}(s)$. In Section 3 it will be proved that the upper bound which was derived in Section 2 can be attained by many $(n,s)$-sequences. This is proved by considering Hamiltonian cycles in a subgraph of $G_{n}(s)$ that contains only the vertices of $(n,s)$-necklaces. By adding $s$ vertices (and $s$ edges) of $G_{n-1}(s)$ from necklaces which contain a cyclic run with $s+1$ zeros a longer acyclic $(n,s)$-sequence is obtained. An efficient algorithm to construct these sequences will be presented. In Section 4 enumeration of the number of $(n,s)$-necklaces is done and a formula for the length of a maximum length $(n,s)$-sequence is derived. In Section 5 a maximum length $(n,s)$-sequence from which the position of each $n$-tuple can be efficiently decoded is constructed. Conclusion, several possible generalizations, and future research are discussed in Section 6.

### v2

@@ Abstract
Free-space quantum key distribution requires to synchronize the transmitted and received signals. A timing and synchronization system for this purpose based on a de Bruijn sequence has been proposed and studied recently for a channel associated with quantum communication that requires reliable synchronization. To avoid a long period of no-pulse in such a system on-off pulses are used to simulate a zero and on-on pulses are used to simulate a one. However, these sequences have high redundancy and low rate. To reduce the redundancy and increase the rate, run-length limited sequences in the de Bruijn graph are proposed for the same purpose. The maximum length of such sequences in the de Bruijn graph is studied and an efficient algorithm to construct a large set of these sequences is presented. Based on known algorithms and enumeration methods, maximum length sequence for which the position of each window can be computed efficiently is presented and an enumeration on the number of such sequences is given.
Index Terms: de Bruijn graph, necklaces, quantum key distribution, rate, RLL sequences.
@@ 1 Introduction
Quantum key distribution are important to prevent quantum computer-based attacks on public key cryptosystems [37]. In a free-space quantum key distribution, one of the important challenges is to synchronize the transmitted and received signals accurately. There are many efforts in designing an efficient and reliable timing and synchronization systems, e.g. [8, 24]. Unfortunately, the suggested systems suffer from a few disadvantages such as slow transmission if for example a clock is used at either end of the transmitter and the receiver. To overcome such problems, in [37] a de Bruijn sequence-based timing and synchronization system is introduced using a beacon with an on-off model. In this model, a sequence of beacon pulses is used to represent a binary de Bruijn sequence. Hence, once a sub-string of beacon pulses is received, its position is also determined uniquely. Furthermore, to consider the timing jitter performance, a long period of no-pulses is forbidden. Assume one pulse slot is used to represent a binary bit. If on-pulse is one and off-pulse is zero, then a long run of zeros in the sequence, which is a long period of no-pulses, would impact the timing jitter. In [37], two pulse slots are suggested to represent a single bit, on-on (i.e., ‘11’) is a one and on-off (i.e., ‘10’) is a zero so that two consecutive no-pulses are avoided. However, this scheme requires $2N$ pulse slots to represent a sequence of length $N$ in the order $n$ de Bruijn graph and it is required to receive a sub-sequence of $2n$ pulse slots to locate its position, i.e., if the sequence is long of about $2N=2^{n+1}$ bits (representing a sequence of length $N$ in the de Bruijn graph), about $2\log N$ pulse slots are required to locate its position instead of $\log N$ pulse slots if any sequence of $N$-pulses was permitted (all logarithms in this paper are in base 2). A second disadvantage in the proposed scheme is that the sequences of consecutive zeros were constrained to length one, while in reality, a few consecutive no-pulses are permitted subject to a constraint that their length will not be larger than a certain threshold $s$. Therefore, there is a target to use less redundant pulse slots to achieve both goals, to synchronize accurately, and to avoid long periods of no-pulses. For this purpose run-length limited (RLL) sequences in the de Bruijn graph are proposed [4, 6], where one pulse is represented by one binary bit. An on-pulse is represented by a one and an off-pulse is represented by a zero. This scheme is more general and more efficient with lower redundancy and higher rate than the ones in the previous work. The scheme combines two concepts, RLL sequences and sequences in the de Bruijn graph which will be defined now. Such sequences are important from engineering point of view. They are formed by using combinatorial properties of the de Bruijn graph. Moreover, the scheme itself is simple and can be applied easily.
@@
RLL sequences are binary sequences in which there is an upper bound on the number of consecutive zeros in a sequence. An $(n,s)$-word is a binary word of length $n$ in which the longest run of consecutive zeros is of length at most $s$. An $(n,s)$-sequence is a sequence whose windows of length $n$ are distinct $(n,s)$-words. The family of sequences in which there are no runs of more than $s$ consecutive zeros was extensively studied due to many applications that require such sequences [22, 23].
@@
The de Bruijn graph of order $n$, $G_{n}$, was defined first by Nicolaas Govert de Bruijn [2] and in parallel by Good [21]. The graph is a directed graph with $2^{n}$ vertices which are represented by the set of all binary words of length $n$. The edges of $G_{n}$ are represented by the $2^{n+1}$ binary words of length $n+1$. There is a directed edge $(x_{0}x_{1}~{}\cdots~{}x_{n-1}x_{n})$ from the vertex $(x_{0}x_{1}~{}\cdots~{}x_{n-1})$ to the vertex $(x_{1}~{}\cdots~{}x_{n-1}x_{n})$. A span $n$ de Bruijn sequence is a cyclic binary sequence in which each binary $n$-tuple is contained exactly once in a window of length $n$. The de Bruijn graph and its sequences associated with the graph were studied extensively [9, 20].
A walk in the graph is a sequence of directed edges such that the end-vertex of one edge is the start-vertex of the next edge. The length of a walk is the number of edges in the walk. A tour is a walk in which the first vertex is also the last one. Any binary sequence can be represented by a walk in $G_{n}$, where any $n$ consecutive symbols represent a vertex and any $n+1$ consecutive symbols represent an edge. A cyclic sequence can be represented by a tour. The consecutive $n$ symbols and $n+1$ symbols represent a walk with its consecutive vertices and consecutive edges, respectively. A path (cycle) in the graph is a walk (tour) with no repeated vertices. In a cycle, each vertex can be considered as the first vertex of the cycle. A trail (circuit) in the graph is a walk (tour) with no repeated edges. A span $n$ de Bruijn sequence can be represented by an Eulerian circuit in $G_{n-1}$, i.e., a circuit which traverses each edge exactly once. In a circuit, the first vertex is also the last one. It can be also represented by a Hamiltonian cycle in $G_{n}$, i.e., a cycle that visit each vertex of $G_{n}$ exactly once. Similarly, any path of length $N$ in $G_{n}$ can be represented by an acyclic sequence of length $N+n-1$ with no repeated $n$-tuples. Each $n$-tuple is associated with a vertex. This sequence can be represented also by a trail in $G_{n-1}$, where each $n$-tuple is associated with an edge. This is demonstrated in Example 1, where also the distinction between cyclic and acyclic sequence is demonstrated. In the rest of the paper if not mentioned, then the sequence is cyclic, but the results are given for both types of sequences.
@@
For $n=3$, the cyclic sequence $[00011101]$ is a span 3 de Bruijn sequence whose cycle is in $G_{3}$ is as follows
$(000)\rightarrow(001)\rightarrow(011)\rightarrow(111)\rightarrow(110)\rightarrow(101)\rightarrow(010)\rightarrow(100)\rightarrow(000).$
In $G_{2}$ this circuit is
[…]
As an acyclic sequence this sequence is written as $0001110100$, i.e., adding two bits (and in general $n-1$ bits) to the cyclic sequence. For the following path in $G_{3}$
$(000)\rightarrow(001)\rightarrow(011)\rightarrow(111)\rightarrow(110)$
its sequence (acyclic) is $0001110$.
[…]
Example 2.
Assume $n=6$ and consider the $(6,2)$-necklace with the vertex $(001011)$. This necklace is of full-order and it contains the six words of length 6 from the cycle
@@
The degenerated necklace $[011]$ of order 6 contains the three vertices in $G_{6}$ which form the cycle of length 3 whose vertices and edges are given by
@@
Now, let
  • $h_{n,s}$ be the number of $(n,s)$-words,
  • $\ell_{n,s}$ be the number of words in all the $(n,s)$-necklaces,
  • $m_{n,s}$ be the maximum length of a cyclic $(n,s)$-sequence.
One of the main results of this contribution is to demonstrate that $m_{n,s}\leq\ell_{n,s}$. The existence of cyclic $(n,s)$-sequences with length $\ell_{n,s}$ are known to exist from the literature with simple constructions [17, 35], and hence we have that $m_{n,s}=\ell_{n,s}$.
In this work, a combination of sequences in the de Bruijn graph that have a run-length constraint on the number of consecutive zeros, is discussed. Another combination of sequences in the de Bruijn graph with local run-length constraint was considered in [5]. An upper bound on the maximum length of cyclic and acyclic $(n,s)$-sequences is proved and sequences which attain this bound are constructed. By their definition, if $s$ is small, these sequences can be used in free-space quantum key distribution with flexibility in the parameters. This is a combinatorial concept motivated by an engineering application problem and can be used in a very simple way for this application.
Let $G_{n}(s)$ be the subgraph of $G_{n}$ induced by all the vertices of $G_{n}$ whose representations are $(n,s)$-words. Our goal is to find the maximum length path and maximum length cycle in $G_{n}(s)$. The definition of $G_{n}(s)$ implies that the edges of this graph also do not have a run of zeros whose length is greater than $s$. Therefore, all the $(n,s)$-words are represented by the edges of $G_{n-1}(s)$. This property will be important in $G_{n-1}(s)$. This implies that a maximum length $(n,s)$-sequence is a trail of maximum length in $G_{n-1}(s)$ and also a path of maximum length in $G_{n}(s)$. For the upper bound on the maximum length of such a sequence, a trail in $G_{n-1}(s)$ (which is a circuit if the sequence is cyclic) will be considered. For the lower bound on the maximum length of such a sequence and constructing many such sequences, paths in $G_{n}(s)$ (which is a cycle if the sequence is cyclic). This is unique as usually it is not required to use different orders of the graph for similar tasks.
@@
The graphs $G_{3}$, $G_{3}(1)$ and $G_{3}(2)$ are depicted in Fig. 1. It is readily verified that a maximum length circuit in $G_{3}(1)$ has length 7 and it is associated with a $(4,1)$-sequence. A maximum length circuit in $G_{3}(2)$ has length 12 and it is associated with a $(4,2)$-sequence. A maximum length cycle in $G_{3}(1)$ has length 4 and it is associated with a $(3,1)$-sequence. A maximum length cycle in $G_{3}(2)$ has length 7 and it is associated with a $(3,2)$-sequence.
Figure 1: The de Bruijn graph $G_{3}$ on the left, $G_{3}(1)$ in the middle and $G_{3}(2)$ on the right.
@@
Note, that a cyclic sequence of length $N$ with no repeated $n$-tuples has $N$ binary digits, $N$ vertices, and $N$ edges in $G_{n}$ and also in $G_{n}(s)$. A related acyclic sequence with no repeated $n$-tuples has $N+n-1$ binary digits, $N$ vertices, and $N-1$ edges in $G_{n}$ and $G_{n}(s)$. Henceforth, unless stated otherwise, all the $(n,s)$-sequences that will be considered are cyclic.
Example 4.
The following cycle in $G_{5}(2)$
$(00110)\rightarrow(01101)\rightarrow(11010)\rightarrow(10101)\rightarrow(01011)\rightarrow(10111)\rightarrow(01111)\rightarrow(11111)$
$\rightarrow(11110)\rightarrow(11101)\rightarrow(11011)\rightarrow(10110)\rightarrow(01100)\rightarrow(11001)\rightarrow(10010)\rightarrow(00101)$
$\rightarrow(01010)\rightarrow(10100)\rightarrow(01001)\rightarrow(10011)\rightarrow(00110)$
forms the $(5,2)$-sequence
$[00110101111101100101]$
of length 20. The sequence contains two words $(00110)$ and $(01100)$ which are two consecutive $(5,2)$-words not in a $(5,2)$-necklace since their necklace contains the words $(00011)$ and $(11000)$, that are not $(5,2)$-words.
$\blacksquare$
@@
The maximum length $(5,2)$-sequence has length 21. Hence, Example 4 indicates that there could be long (almost the same length as the maximum length or maybe of maximum length) $(n,s)$-sequences which contain $(n,s)$-words that are not contained in $(n,s)$-necklaces. The possibility of such maximum length $(n,s)$-sequences with $(n,s)$-words which are not contained in $(n,s)$-necklaces cannot be ruled out by simple verification. This implies that it is not straightforward to know the maximum length of an $(n,s)$-sequence. Moreover, the length of a maximum length acyclic $(n,s)$-sequence is not obtained from the maximum length $(n,s)$-sequence by adding $n-1$ bits (as done for a de Bruijn sequence), since more $(n,s)$-words can be added to the cyclic sequence.
Observing that RLL sequences can be applied efficiently for quantum key distribution is the first contribution of this paper from an engineering point of view. This is the motivation for this paper. The main contributions of this work can be summarized as follows:
  1. 1.
    A proof for an upper bound on the length of the constructed cyclic and acyclic sequences, i.e., it is proved that $m_{n,s}\leq\ell_{n,s}$. It was known before that there are $(n,s)$-sequences of length $\ell_{n,s}$, but it was not known that these are the $(n,s)$-sequences of maximum length. Our proof implies that $m_{n,s}=\ell_{n,s}$. The proof also implies that a maximum length cyclic $(n,s)$-sequence contains exactly all the words of all the $(n,s)$-necklaces. Related results are obtained for acyclic $(n,s)$-sequences.
  2. 2.
    An efficient construction for a large set of such sequences using a tailored designed storage that satisfies all the requirements of $(n,s)$-sequences. Constructions for one $(n,s)$-sequence was known before, but these constructions cannot be adapted for constructing a large set of $(n,s)$-sequences.
The contributions of the paper are presented in Sections 2 and 3. Sections 4 and 5 are results which are either straightforward consequences or were obtained in previous publications. These results are presented for completeness. For simplicity we usually assume that $s<n-1$ since if $s\geq n$ then an span $n$ de Bruijn sequence is an $(n,s)$-sequence and if $s=n-1$, then a shortened span $n$ de Bruijn sequence (one zero is removed from the run of $n$ consecutive zeros) is an $(n,s)$-sequence.
@@
The rest of the paper is organized as follows. In Section 2 we consider an upper bound on the maximum length of $(n,s)$-sequences which are cycles in $G_{n}$ (and also in $G_{n}(s)$). We prove that this bound is equal to the number of words in the $(n,s)$-necklaces, i.e., $m_{n,s}\leq\ell_{n,s}$ and it can be attained only by the vertices of these necklaces. The trail in $G_{n-1}(s)$ with the maximum length can have another $s$ edges (and $s$ vertices) from $G_{n-1}(s)$. The detailed proof is based on the necklaces obtained by the edges of $G_{n-1}(s)$. In Section 3 it will be proved that the upper bound which was derived in Section 2 can be attained by many $(n,s)$-sequences which can be constructed efficiently. This is proved by considering Hamiltonian cycles in a subgraph of $G_{n}(s)$ that contains only the vertices of the $(n,s)$-necklaces. By adding $s$ vertices (and $s$ edges) of $G_{n-1}(s)$ from necklaces which contain a cyclic run with $s+1$ zeros a longer acyclic $(n,s)$-sequence is obtained. An efficient algorithm to construct these sequences will be presented. Although the type of algorithm which is presented is not new, the choice of keys to form a very large set of such sequences is new. In Section 4 enumeration of the number of $(n,s)$-necklaces is done and a formula for the length of a maximum length $(n,s)$-sequence is derived. In Section 5 a maximum length $(n,s)$-sequence from which the position of each $n$-tuple can be efficiently decoded is constructed. The construction is an adaptation of a well-known method. The section considers the related literature on these types of constructions. Conclusion, several possible generalizations, and future research are discussed in Section 6.